OpenStudy (anonymous):

A child lies on the ground and looks up at the top of a 17-ft tree nearby. The child is 14 ft away from the tree. What is the angle of elevation from the child to the top of the tree? Round to the nearest whole degree.

OpenStudy (anonymous):

\[180 \left.\tan ^{-1}\left(\frac{17}{14}\right)\right/\pi =51{}^{\circ} \]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i gave you a medal

OpenStudy (anonymous):

Thank you for the medal.